# SoMeR: Multi-View User Representation Learning for Social Media

### v1

@@ 4 Detecting Drivers of Information Operations
Supervised Classification
We use a two-step approach to classify IO drivers. First, we pre-train our model in the self-supervised manner with both IO driver and control users using all of their posts and meta-data, but without any IO label. For textual post preprocessing, we remove user mentions, URLs, emojis and all non-ASCII characters, but retain the hashtags. This is a multi-lingual datasets including more than 50 languages. Therefore we compute the sentence-BERT embedding for each post using the stsb-xlm-r-multilingual [^1] https://huggingface.co/sentence-transformers/stsb-xlm-r-multilingual and compute the first five components from PCA as its textual features. This is to reduce the number of triplets and to reduce computation complexity. We aggregate data for each user by summing up their PCA embeddings of their posts in 3-day intervals and taking the middle day in the interval as the corresponding timestamp. Other intervals can be used, although this value acts as a reasonable balance for low-activity accounts. This gives us the temporal and textual features. For profile features, we use number of followings and followers, as accounts such as news media outlets with lots of followers can behave very differently from other types users. For network link prediction part, we use the repost network (i.e.retweet network) because co-repost has been shown as a potential indicator of an IO driver [28]. With all these features, we train the model to learn a user embeddings space in which users with similar behaviors are closer together (cf. Figure 1).
In the next step, we perform supervised fine-tuning on the learned model parameters with an additional two-layer feed-forward network for binary classification. This network has a linear layers with a hidden dimension of 128, a ReLU activation, a dropout layer with rate of 0.3, a batch normalization, and a second linear layer that project embeddings onto $\mathbb{R}$ for binary prediction. The sigmoid function is then applied and a binary cross-entropy function is used as the loss. We use same hyperparameters described in § 3.
@@
We compare our method with predominantly used IO driver detection methods [28]. Based on the observation that IO drivers have abnormally similar behaviors, this method first construct similarity networks among users based on various features such as co-sharing, and identify users with similarity in the top-percentile as IO drivers. Following previous works [28, 23], we detect IO drivers based on (1) co-repost and (2) co-URL sharing behaviors if cosine similarity is above or at the 99.5-th percentile, (3) hashtag sequence (using a minimum sequence of 5 identical hashtags in the same order within a post) and (4) text similarity in averaged BERT embeddings over all posts from a user, with a cosine similarity threshold of 0.7.
Table 2: Model Performance on Detecting Information Operation Drivers.
  F1-Scores ROC-AUC
China Egypt-UAE Venezuela China Egypt-UAE Venezuela
@@
Co-Repost 0.00 0.15 0.26 0.50 0.54 0.58
Co-URL 0.19 0.27 0.30 0.45 0.74 0.52
Hashtag-sequence 0.08 0.20 0.05 0.51 0.56 0.51
Text Similarity 0.13 0.93 0.82 0.53 0.94 0.85
@@
Temporal $0.96\pm 0.01$ $0.41\pm 0.13$ $0.57\pm 0.06$ $0.99\pm 0.01$ $0.95\pm 0.01$ $0.90\pm 0.02$
Textual $0.97\pm 0.01$ $0.69\pm 0.06$ $0.67\pm 0.04$ $0.99\pm 0.01$ $0.96\pm 0.02$ $0.96\pm 0.01$
Temporal+Textual $0.98\pm 0.01$ $0.77\pm 0.05$ $0.77\pm 0.04$ $0.99\pm 0.01$ $0.98\pm 0.01$ $0.97\pm 0.01$
SoMeR $\mathbf{0.99\pm 0.01}$ $0.85\pm 0.04$ $\mathbf{0.82\pm 0.04}$ $\mathbf{0.99\pm 0.01}$ $\mathbf{0.99\pm 0.01}$ $\mathbf{0.97\pm 0.01}$
Model Performance
We demonstrate how our framework can effectively learn from user behaviors and network connections, and adequately detect IO drivers. Table 2 show the outstanding performance of our method from the F1-scores and ROC-AUC. Our models are evaluated from 10 random data splits. First we compare our method with the baselines. For all three campaigns, we outperform the co-repost, co-URL and hashtag-sequence baselines. Compared to the text similarity baseline, our method outperforms on China and Venezuela campaigns, and has a higher ROC-AUC score but a lower F1-score on the Egypt-UAE campaign. We hypothesize that our method utilizing only the first five PCA components on post BERT embeddings versus the text similarity baseline using the full BERT embeddings might result in this gap. This can be addressed by using the full BERT embeddings when having more powerful GPUs. Nonetheless, the high F1-scores and AUC scores we achieve demonstrate the effectiveness of our method.
The performance of different baselines also reveal some interesting facts. Co-sharing and hashtag usage seem to be less distinguishable from non-IO accounts in all three campaigns. On the other hand, by just using text similarity can achieve a very high performance on Egypt-UAE and Venezuela datasets, but the text similarity baseline has relatively low performance on the China dataset, indicating that different campaigns utilize different tactics. These observations demonstrate that only by looking at one type of suspicious behavior pattern is not enough when detecting IO drivers.
Ablation
@@
Next, we perform an ablation study on our model to dissect the impact of different features and modules we use. The Temporal model only uses timestamps and three-day post counts as the feature but do not use any textual or network features. The Textual model only uses timestamps and the average textual embedding over three-day intervals, which does not reflect temporal activity. The Temporal+Textual model uses timestamps and the summed textual embeddings over three-day intervals, as described in § 4 Supervised Classification. This reflects both temporal activity and textual features. The full model SoMeR incorporates network feature on top of the temporal and textual features. We observe that the full model has better performance than any of the ablated models, implying the benefit of including the network link prediction objective. In addition, Temporal+Textual model performs better than Temporal and Textual models. These indicate that each of the four features plays a role to detect IO drivers.
5 Uncovering Shifts in Polarized Discussions
@@
The U.S. society has grown increasingly more polarize. Not only do liberals and conservatives hold sharply different opinions on a range of issues [3], but they also have more negative feelings towards members of the other party, compared to members of their own party [19]. These differences show up not only in political speech, but also in the everyday behaviors [10], on on social media platforms, where liberals and conservatives segregate themselves in different echo chambers, they are reflected in the network structure [6]. Prior research has found that events can shift public opinion, polarizing the population [16, 20, 32]. While many alternative methods for measuring polarization on social media exist, e.g., using network analysis [6], interaction behaviors and emotions [9], and community embeddings [41], they are not well suited for measuring shifts in polarization. Leveraging our multi-view user representation learning, we provide a new way to measure changes in polarization after significant events by tracking user embeddings, and apply it to measure polarization in the online discussions about abortion, a highly contentious issue in the American society. Our framework allows us to isolate specific topics and identify ones that grew more polarized.
@@
On June 24, 2022, SCOTUS struck down federal protections for abortion rights. This event sparked massive online discussions, in which users with different political ideologies expressed distinct views [32]. We study a public dataset [4] containing tweets with abortion-related keywords, such as “roevswade”, “prochoice”, and “prolife”. The data spans the entire year of 2022. We select English posts in the U.S. from users with at least 20 posts. This gives us about 10 million posts from 0.1 million users. Using methodology described in Rao et al. [32], we identify each user’s political ideology (liberal/conservative), leaving us with 103K liberals and 18K conservatives. We use the learned user embeddings to track how these two ideological populations change in the user embedding space.
@@
We start with pre-training, which learns a user representation space for all liberal and conservative users using their posting histories in 2022. We perform the same text preprocessing as described in § 4, except that the sentence-BERT model we use for this English-only dataset is sentence-transformers/all-mpnet-base-v2222https://huggingface.co/sentence-transformers/all-mpnet-base-v2. We use following and follower counts as profile features, and use a repost (i.e., retweet) network to train link prediction objective. Figure 2(a) shows a t-SNE [39] representation of the learned embedding space for all users for the year 2022. We see the separation of populations, with conservatives clustered in some regions, whereas liberals, who are 85% of all data, are distributed across the entire space.
Figure 2: (a) t-SNE [39] of user embedding space learned with both ideology populations over the entire year of 2022. (b) t-SNE of the space in (a) is fine-tuned with data from elite users with known ideologies, making the embedding space more politically-aware.
Next, we fine-tune the model via few-shot learning with political elites. [^3] https://github.com/sdmccabe/new-tweetscores?tab=readme-ov-file These correspond to well known people, e.g., U.S. politicians, with a known ideology. We found 358 in our abortion data. This fine-tuning step trains the model to predict user ideology, thereby aligning the embedding space to political ideology. For the analysis we discuss below, we test in both the original and the fine-tuned politically-aware embedding and find the same trends. Figure 2(b) shows a t-SNE visualization of the politically-aware user embeddings after few-shot fine-tuning. We see that the conservative users have moved closer in one direction, indicating the effectiveness of fine-tuning at distinguishing these disparate groups.
In this politically-aware embedding space, to identify shifts in ideological polarization driven by the SCOTUS ruling, we measure how embeddings changed for the same users after the ruling. We first establish a baseline period when users were not affected. There was a leak about this ruling on May 3rd, 2022. Therefore we select the period of January 1st to May 2nd 2022 as the baseline period to avoid interference from this event. We then determine the period to observe the impact of ruling to be June 24th to November 11th 2022. We select this end date to reduce the confounding effect of the 2022 US midterm elections. Next, we select users who posted in both time periods, take their posting history in these two periods separately, and project these users onto the same politically-aware embedding space we have learned. By visualizing the embeddings of the same users in the baseline period and after the SCOTUS ruling, we find a clear shift especially in the conservative population (Figure 3). In the baseline period, conservative users were more uniformly distributed in the t-SNE embedding space, but they moved closer together after the SCOTUS ruling. This indicates that the conservative users became more similar in the content of their posts or in their behaviors.
@@
To dig deeper, we explored how this effect depends on the topic users discuss. Rao et al. [32] identified topics discussed in each post, including religion, bodily autonomy, fetal rights and women’s health. We created one subset with posts related to liberal-centric topics, i.e. bodily autonomy and women’s health, and another subset with posts related to conservative-centric topics, like religion and fetal rights. Then we perform the same analysis for each subset. Figure 4 shows a consistent overall trend that users with same ideology move closer and users with different ideologies move farther away. Interestingly, we also observe that each population coalesced when discussing partisan topics promoted by the opposite ideology - in-group neighbors of conservative users increased more on Liberal-centric topics than on Conservative-centric topics, and similarly in-group neighbors of liberals increased more on Conservative-centric topics than on Liberal-centric topics. Users “united against a common enemy” in these online discourses.
@@
6 Conclusion

### v2

@@ Methods
Pre-training and Supervised Classification
We use a two-step approach to classify IO drivers. First, we pre-train our model in the self-supervised manner with both IO driver and control users using all of their posts and meta-data, but without any IO label. For textual post preprocessing, we remove user mentions, URLs, emojis and all non-ASCII characters, but retain the hashtags. This is a multi-lingual datasets including more than 50 languages. Therefore we compute the sentence-BERT embedding for each post using the stsb-xlm-r-multilingual  [^1] https://huggingface.co/sentence-transformers/stsb-xlm-r-multilingual, and compute the first five components from PCA as its textual features. Due to GPU limit, we use this technique to reduce the number of triplets and computation complexity. We aggregate data for each user by summing up their PCA embeddings of their posts in 3-day intervals and taking the middle day in the interval as the corresponding timestamp. We have conducted experiments with 1-day, 3-day and 5-day window sizes, and choose to use 3-day as it keeps the data relatively fine-grained and also aggregate data well for low-activity accounts. This gives us the temporal and textual features. For profile features, we use number of followings and followers, as accounts such as news media outlets with lots of followers can behave very differently from other types users. For network link prediction part, we use the repost network (i.e. retweet network) because co-repost has been shown as a potential indicator of an IO driver (Pacheco et al. 2021). With all these features, we train the model to learn a user embeddings space in which users with similar behaviors are closer.
In the next step, we perform supervised fine-tuning on the learned model parameters with an additional two-layer feed-forward network for binary classification. This network has a linear layers with a hidden dimension of 128, a ReLU activation, a dropout layer with rate of 0.3, a batch normalization, and a second linear layer that project embeddings onto $\mathbb{R}$ for binary prediction. The sigmoid function is then applied and a binary cross-entropy function is used as the loss. We use same hyperparameters described in Methods.
Model Performance and Ablation
@@
We compare our method with common IO driver detection methods (Pacheco et al. 2021). Based on the observation that IO drivers have abnormally similar behaviors, these methods identify users with unusual similarity based on (1) co-repost and (2) co-URL sharing behaviors if cosine similarity is above or at the 99.5-th percentile, (3) hashtag sequence (using a minimum sequence of 5 identical hashtags in the same order within a post) and (4) text similarity in averaged BERT embeddings over all posts from a user, with a cosine similarity threshold of 0.7 (parameters based on prior works (Pacheco et al. 2021; Luceri et al. 2024).)
In addition, we also compare with a strong multi-view URL baseline SATAR (Feng et al. 2021). This approach combines textual, profile and network features, and the authors uses it for social bot detection. However, the major differences between SATAR and our framework are (1) SATAR does not utilize temporal features; (2) SATAR uses follower count prediction as the self-supervised objective, which is suitable for bot detection but less generalizable for other use cases, whereas SoMer uses both contrastive learning and network prediction objectives for self-supervision signals; (2) SATAR uses the Long Short-Term Memory model (LSTM) (Graves and Graves 2012) with word2vec embeddings (Church 2017) whereas SoMer uses more advanced transformer module to learn contextual information from user posting histories. The details of evaluating SATAR are included in Appendix.
Performance
Table 2 shows the performance of different methods. Our models are evaluated on 10 random training-validation-test data splits. First we compare our method with the baselines. For China and Venezuela campaigns, we outperform or match all the baselines. For the gap between the text similarity F1-score and ours for the Egypt-UAE campaign, we hypothesize that it is due to our method utilizing only the first five PCA components on post BERT embeddings, versus the text similarity baseline using the full BERT embeddings. In addition, Egypt-UAE is the smallest campaign, and pre-training on a small dataset can also result in less ideal performance (see more analysis in Appendix). Moreover, our full model outperforms SATAR in all scenarios. We observe that SATAR’s pre-training step is ineffective, with the fine-tuning step contributing most significantly to its overall performance. SATAR’s self-supervised pre-training objective, which predicts user follower counts, easily leads to overfitting. For example, for the Chinese campaign, F1-score for predicting follower counts reaches 0.88 but the F1 for predicting IO drivers is only 0.13 during pre-training. After fine-tuning, the F1 for predicting IO drivers reaches 0.98 as in Table 2. This reliance on labeled data highlights SATAR’s limited generalizability to scenarios with sparse labels. In conclusion, the consistently high F1-scores SoMer achieves demonstrate the effectiveness of our method.
Table 2: F1-Scores on Detecting IO Drivers.
  China Egypt-UAE Venezuela
@@
Co-Repost 0.00 0.15 0.26
Co-URL 0.19 0.27 0.30
Hashtag-sequence 0.08 0.20 0.05
Text Similarity 0.13 0.93 0.82
SATAR $0.98\pm 0.01$ $0.72\pm 0.03$ $0.61\pm 0.04$
@@
Temporal $0.96\pm 0.01$ $0.41\pm 0.13$ $0.57\pm 0.06$
Textual $0.97\pm 0.01$ $0.69\pm 0.06$ $0.67\pm 0.04$
Temporal+Textual $0.98\pm 0.01$ $0.77\pm 0.05$ $0.77\pm 0.04$
SoMeR $\mathbf{0.99\pm 0.01}$ $0.85\pm 0.04$ $\mathbf{0.82\pm 0.04}$
Ablations
@@
Next, we perform an ablation study on our model to dissect the impact of different features and modules we use. The Temporal model only uses timestamps and three-day post counts as the feature but do not use any textual or network features. The Textual model only uses the average textual embedding over three-day intervals, which does not reflect temporal activity. The Temporal+Textual model uses timestamps and the summed textual embeddings over three-day intervals, as described in Detecting IO Drivers. This reflects both temporal activity and textual features. The full model SoMeR incorporates network feature on top of the temporal and textual features. All these models include the profile features. We observe that the full model has better performance than any of the ablated models, implying the benefit of including the network link prediction objective. In addition, Temporal+Textual model performs better than Temporal and Textual models alone, indicating that each of the four features plays a role to detect IO drivers. For the China campaign, all ablated and full models have high F1 scores. First, this implies the importance of the temporal features. The Textual model uses average textual embeddings over three-day intervals. This lowers but does not completely eliminate temporal factors, and can result in the similar behaviors to the Temporal model. We further confirm the importance of combining different modules by t-SNE of the embeddings generated by different modules in Figure 6 in Appendix.
The performance of different SoMer modules and different baseline approaches vary across different campaigns. For example, the baseline Text Similarity performs well for Egypt-UAE and Venezuela but poorly for China. On the other side, our models using Temporal features show excellent performance on the China campaign. This shows that campaigns use different tactics and highlights that relying on a single behavior pattern is insufficient, whereas our framework’s multi-view approach enables better detection.
Uncovering Shifts in Polarized Discussions
@@
The U.S. society has grown increasingly more polarize. Not only do liberals and conservatives hold sharply different opinions on a range of issues (Center 2017), but they also have more negative feelings towards members of the other party, compared to members of their own party (Iyengar and Krupenkin 2018). These differences show up not only in political speech, but also in the everyday behaviors (DellaPosta et al. 2015), on on social media platforms, where liberals and conservatives segregate themselves in different echo chambers, they are reflected in the network structure (Conover et al. 2011). Prior research has found that events can shift public opinion, polarizing the population (Hebbelstrup Rye Rasmussen and Petersen 2023; Jiang et al. 2020; Rao et al. 2023). While many alternative methods for measuring polarization on social media exist, e.g., using network analysis (Conover et al. 2011), interaction behaviors and emotions (Del Vicario et al. 2016), and community embeddings (Waller and Anderson 2021), they are not well suited for measuring shifts in polarization. Leveraging our multi-view user representation learning, we provide a new way to measure changes in polarization after significant events by tracking user embeddings, and apply it to measure polarization in the online discussions about abortion, a highly contentious issue in the American society. Our framework allows us to isolate specific topics and identify ones that grew more polarized.
@@
On June 24, 2022, the SCOTUS struck down federal protections for abortion rights. This event sparked many online discussions, in which users with different political ideologies expressed distinct views (Rao et al. 2023). We study a public dataset (Chang et al. 2023) containing tweets with abortion-related keywords, such as “roevswade”, “prochoice”, and “prolife”. The data spans the entire year of 2022. We select English posts in the U.S. from users with at least 20 posts. This gives us about 10M posts from 121K users. Using methodology described in Rao et al. (2023), we identify each user’s political ideology (liberal/conservative), leaving us with 103K liberals and 18K conservatives. We use the learned user embeddings to track how these two ideological populations change in the user embedding space.
@@
We start with pre-training, which learns a user representation space for all liberal and conservative users using their posting histories in 2022. We perform the same text preprocessing as described in Detecting IO Drivers, except that we use sentence-transformers/all-mpnet-base-v2222https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for this English-only dataset. We also use following and follower counts as profile features and use a repost (i.e., retweet) network to train link prediction objective.
Next, we fine-tune the model via few-shot learning with 358 political elites [^3] https://github.com/sdmccabe/new-tweetscores, such as U.S. politicians, with a known ideology. This aligns the embedding space to political ideology. We use a two-layer feed-forward network on top of SoMer and fine-tune the model as described in Detecting IO Drivers. Figure 7 in Appendix shows t-SNE (Van der Maaten and Hinton 2008) representations of the learned embedding space for all users for the year 2022 both with and without fine-tuning. We see the separation of populations, with conservatives clustered in some regions, whereas liberals, who are 85% of all data, are distributed across the entire space. The analysis we discuss below describes results using the fine-tuned politically-aware embedding, but we find the same trends using the pre-trained embeddings.
In this politically-aware embedding space, to identify shifts in ideological polarization driven by the SCOTUS ruling, we measure how embeddings changed for users before versus after the ruling. We select the period of January 1st to May 2nd 2022 as the baseline period to avoid interference from a leak about this ruling on May 3rd, 2022. We then determine the period to observe the impact of ruling to be June 24th to November 11th 2022. We select this end date to reduce the confounding effect of the 2022 US midterm elections. Next, we select users who posted in both time periods, take their posting history in these two periods separately, and project these users onto the same politically-aware embedding space we have learned. By visualizing the embeddings of the same users in the baseline period and after the SCOTUS ruling, we find a clear shift especially in the conservative population (Figure 2). In the baseline period, conservative users were more uniformly distributed in the t-SNE embedding space, but they moved closer together after the SCOTUS ruling. This indicates that the conservative users became more similar in the content of their posts or in their behaviors.
@@
To dig deeper, we explored how this effect depends on the topic users discuss. Rao et al. (2023) identified topics discussed in each post, including religion, bodily autonomy, fetal rights and women’s health. We created one subset with posts related to liberal-centric topics, i.e. bodily autonomy and women’s health, and another subset with posts related to conservative-centric topics, like religion and fetal rights. Then we perform the same analysis for each subset. Figure 3 shows a consistent overall trend that users with same ideology move closer and users with different ideologies move farther away. Interestingly, we also observe that each population coalesced when discussing partisan topics promoted by the opposite ideology - in-group neighbors of conservative users increased more on Liberal-centric topics than on Conservative-centric topics, and similarly in-group neighbors of liberals increased more on Conservative-centric topics than on Liberal-centric topics. Users “united against a common enemy” in these online discourses.
@@
Predicting Participation in Hate Subreddits
Online hate communities have numerous harmful effects on users(Schmitz et al. 2024) by radicalizing them (START 2017) and increasing hate crimes (FBI 2021). While there has been interest in what drives people to hate groups (Russo et al. 2024), this analysis was often based on case studies of individual communities. Less-understood is why users may be susceptible to extremism and hate. We aim to use SoMeR to predict whether a Reddit user who had never participated in any hate subreddits will become active in one weeks or even months in the future (on average a user’s last post fed into SoMeR was 22 days before their first post in a hate subreddit). We believe this can help explain what drives people to join hate groups.
Data
We use a list of 168 hate subreddits collected by (Hickey et al. 2024), which consists of large banned subreddits with hate speech. We manually inspected the subreddits to verify that they were indeed hate subreddits. We then extract all users who have ever posted in these subreddits until June, 2022 using data stored in Academic Torrent  [^4] https://files.pushshift.io/reddit/. We take a random 200K subsample of all the hate users we collected. To construct a control user group, we collect the user names of 100K random users who ever made a post on Reddit until June, 2022.
We then collect all posts made by each of these random and hate users, and matched random and hate users such that random users and hate users were active at roughly the same time to make them less trivially distinguishable (Luceri et al. 2024). To match random users with hate users, we kept track of which hate users had already been paired. For each random user, we filtered hate users who hadn’t been matched, ensuring their first post was before becoming active in a hate subreddit. Then, we randomly selected a hate user active during the same time period as the random user. During training and testing, for each hate user, we only use their posts and comments before the date they became active in a hate subreddit, and we use this same time range for each matched control user. Within this dataset, we randomly select 4K matched pairs (8K users total) to have a manageable data size for fast experimenting. We keep the overall time range of the dataset to be January 1, 2013 to December 31, 2022, because of the low amount of data before 2013. We then remove low-activity users who have posted and commented less than 10 times. This results in 3771 hateful users and 3444 control users, and 2.3 million posts and comments in total. We split data by 70%-15%-15% for training, validation and testing, respectively.
Pre-training and Supervised Classification
We perform pre-training and supervised fine-tuning to predict whether a user would become active in a hate subreddit later. We first use the same data processing procedure as in Pre-training and Supervised Classification, pre-processing text by removing URLs, emojis and all non-ASCII characters, using sentence-BERT all-mpnet-base-v2 to compute the embeddings , and computing the first five components from PCA as the textual features. We aggregate data for each user by summing up their PCA embeddings of their posts in 3-day intervals and taking the middle day in the interval as the corresponding timestamp. This provides us with both the temporal and textual features. For the network module, a reasonable option we have attempted is to use co-subreddit membership as the linkage (similar users participate in the same subreddits). However, users join subreddits at different times, resulting in different co-activity network from time to time. These co-activity networks for user A may include user B who already joined the hate group, creating information leakage between training and testing. Moreover, because there is no fixed co-activity network, incorporating this feature into the training is very data-intensive and computationally infeasible. For this reason, we had to exclude the co-subreddit membership network feature, but encode users’ subreddit membership information as the profile feature, by gathering the 50 most-posted subreddits of each user, and use a lookup-table encoder to generate the profile embeddings.
With all the above features, we then pre-train the model to learn a user embeddings space in which users with similar behaviors are closer together. Next, we perform supervised fine-tuning using a two-layer feed-forward network for binary prediction of whether the user will become active in a hate subreddit later (hyperparameters and training procedure same as in Detecting IO Drivers).
Model Performance and Ablation
There is no prior task-specific research for predicting user participation in hate communities, hence we first compare SoMeR with a simple baseline—using sentence-BERT embeddings averaged over user posts prior to becoming active in a hate subreddit to train the same two-layer feed-forward network for binary prediction. We also evaluate the multi-view user representation model SATAR as a strong baseline (see SATAR evaluation details in Appendix.) We evaluate all models with randomly bootstrapping training, validation and test data 10 times. Table 3 shows that SoMeR significantly outperforms the BERT baseline by 9% and SATAR by 20% on F1-scores, indicating the effectiveness of our method. SATAR has surprisingly lower performance, possibly because its self-supervised objective is not fitted for detecting hate behavior.
Table 3: F1 Scores on Predicting User Participation in Hate Subreddits.
  F1-Scores
BASELINE
BERT $0.69\pm 0.01$
SATAR $0.59\pm 0.05$
OURS
Temporal $0.74\pm 0.01$
Textual $0.76\pm 0.01$
Temporal+Textual $0.77\pm 0.00$
Temporal+Textual+Profile $\mathbf{0.78\pm 0.01}$
Table 3 also shows our ablation study, where we see that the Temporal+Textual model outperforms Temporal only and Textual only models, and using Temporal, Textual and Profile further improves the F1 score. This indicates that all these modules are important. We confirm this with t-SNE plots of embeddings generated by different ablated models in Appendix Figure 8.
Conclusion